### v1

# InstantViR: Real-Time Video Inverse Problem Solver with Distilled Diffusion Prior

Abstract
[…]
footnotetext: Corresponding authors.
1 Introduction
Videos have become the dominant medium for communication, entertainment, and information in the digital age. Consequently, the task of solving video inverse problems—reconstructing a high-quality, clean video $\boldsymbol{x}$ from degraded measurements $\boldsymbol{y}$—is of pivotal importance, with applications ranging from enhancing consumer footage and restoring historical films to improving medical imaging and autonomous driving systems [chang2025warped, daras2024warped, kwon2025solving, kwon2024visionxlhighdefinitionvideo, zhang2025step, zou2025flair]. This reconstruction task is inherently ill-posed, as a single measurement $\boldsymbol{y}$ can correspond to many possible clean videos $\boldsymbol{x}$. To solve this ambiguity, a standard approach is to cast video reconstruction as a Bayesian inverse problem, where the goal is to sample from the posterior distribution $p(\boldsymbol{x}|\boldsymbol{y})$:
$p(\boldsymbol{x}|\boldsymbol{y})\propto p(\boldsymbol{y}|\boldsymbol{x})p(\boldsymbol{x}).$ (1)
This formulation elegantly decouples the problem into two components: (1) the likelihood $p(\boldsymbol{y}|\boldsymbol{x})$, governed by the physics of the degradation (e.g., blur, noise, or masking); and (2) the prior $p(\boldsymbol{x})$, which captures the high-dimensional spatiotemporal statistics of natural videos. The central challenge in video inverse problems is therefore how to represent and compute with this video prior $p(\boldsymbol{x})$ in a way that is both expressive and computationally tractable.
[…]
Recently, diffusion models [ho2020denoising, song2020score, sohl2015deep] have become the dominant tool for learning powerful priors, yielding state-of-the-art results for image inverse problems [chung2022diffusion, mardani2023variational, rout2023solving, bai2024blind, lee2024diffusion]. A natural first step for videos is to repurpose pre-trained image diffusion models [ho2020denoising, rombach2022high] for video restoration. Methods along this line [daras2024warped, kwon2025solving, kwon2024visionxlhighdefinitionvideo, zhang2025step] process videos frame-by-frame and enforce consistency only through empirical temporal regularizers (e.g., optical flow constraints or batched noise). While convenient, this image-native strategy has two key drawbacks: (1) it uses priors that ignore rich spatiotemporal dynamics, often leading to flicker and temporal inconsistency; and (2) it still relies on slow, iterative sampling to solve Eq. 1, which is ill-suited for real-time deployment.
A more principled alternative is to adopt native video diffusion models [wan2.1, wan2025], which are trained directly on videos and provide much stronger temporal priors, producing visually coherent sequences. However, when used for inverse problems, these models inherit and amplify the computational burden of diffusion: posterior sampling [chung2022diffusion] in Eq. 1 requires long iterative trajectories in a much higher-dimensional video space, making inference even more expensive. As a result, current diffusion-based approaches face a stark trade-off: either use image diffusion priors that are temporally weak and slow, or video diffusion priors that are temporally strong but prohibitively slower and incompatible with real-time, streaming applications.
In this work, we argue that this trade-off is not fundamental. We show that one can retain the rich, temporally consistent prior of a strong video diffusion model without inheriting its prohibitive sampling cost. To this end, we introduce InstantViR, a variational framework that recasts video reconstruction as amortized inference via video diffusion distillation [luo2023diff, huang2024flow, wang2024integrating, yin2024one, yin2024slow, zhou2024score, lee2024diffusion]. Instead of solving a separate, slow optimization for $p(\boldsymbol{x}|\boldsymbol{y})$ at test time, InstantViR learns a universal one-step solver that maps a degraded video $\boldsymbol{y}$ directly to its reconstruction $\boldsymbol{x}$.
Our method uses an asymmetric teacher–student design. A slow bidirectional video diffusion model (e.g., Wan2.1 [wan2.1]) serves as the teacher, which together with the measurement operator (blur, masking, downsampling, text-guided edits, etc.) defines the target posterior. A fast causal autoregressive student is then trained to approximate this posterior in a single forward pass. Training is paired-data-free: conditioned on the degraded measurement $\boldsymbol{y}$, we align the reconstruction with the teacher’s prior while enforcing physics-based data fidelity, yielding a self-supervised approximation to the video posterior. As a result, the same framework seamlessly adapts to streaming random inpainting, Gaussian deblurring, super-resolution, and language-guided restoration/editing by simply changing the measurement model and the conditioning prompt of the video diffusion prior.
To make the system truly real-time, we also tackle the often-overlooked VAE bottleneck. We introduce a teacher-space regularized latent adaptation scheme that allows us to replace the original video VAE with LeanVAE [cheng2025leanvae]—an ultra-efficient spatiotemporal tokenizer with a lightweight Neighborhood-Aware Feedforward backbone and wavelet-based channel compression—while keeping the distilled solver consistent with the teacher prior. Combined with our one-step causal solver, this design allows InstantViR to deliver diffusion-level reconstructions with up to 100× speedup over iterative video diffusion priors, while maintaining high fidelity and temporal coherence.
Concretely, InstantViR contributes: (i) a paired-data-free variational distillation framework that converts native video diffusion priors into one-step amortized solvers for diverse video inverse problems; (ii) a streaming causal inverse architecture with block-wise attention and KV caching for low-latency, high-fidelity reconstruction; and (iii) a teacher-space regularized LeanVAE integration that overcomes latent distribution shift and, to our knowledge, yields the first real-time, interactive pipeline for streaming video reconstruction with diffusion-level quality at over 35 FPS.
2 Related Works
2.1 Video Diffusion Models as Priors
Diffusion models [ho2020denoising, song2020score, sohl2015deep] have emerged as the state-of-the-art to serve as powerful priors $p(\boldsymbol{x})$ that can accurately model the complex statistics of natural videos. These models define the prior $p(\boldsymbol{x})$ through a learned reverse-time stochastic differential equation (SDE) that denoises a simple Gaussian distribution $p(\boldsymbol{x}_{T})$ back to the data distribution $p(\boldsymbol{x}_{0})$. Concretely, for a clean video $\boldsymbol{x}_{0}$, a fixed forward process defines how noise is added:
$d\boldsymbol{x}_{t}=f(t)\boldsymbol{x}_{t}\,dt+g(t)\,d\mathbf{w}_{t},$ (2)
where $t\in[0,T]$, $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $\mathbf{w}_{t}$ is a standard Wiener process. The generative prior $p(\boldsymbol{x})$ is defined by the corresponding reverse-time SDE:
$d\boldsymbol{x}_{t}=\bigl[f(t)\boldsymbol{x}_{t}-g(t)^{2}\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})\bigr]\,dt+g(t)\,d\bar{\mathbf{w}}_{t},$ (3)
where the key component, the score function $\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})$, is approximated by a neural network $s_{\theta}(\boldsymbol{x}_{t},t)$. This network $s_{\theta}$ is trained by optimizing the denoising score matching (DSM) objective [song2019generative] with $\boldsymbol{x}_{t}=\alpha_{t}\boldsymbol{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$:
$\mathcal{L}_{\mathrm{DSM}}=\mathbb{E}_{t,\boldsymbol{x}_{0},\boldsymbol{\epsilon}}\!\Bigl[\lambda(t)\,\bigl\|s_{\theta}(\boldsymbol{x}_{t},t)+\tfrac{\boldsymbol{\epsilon}}{\sigma_{t}}\bigr\|^{2}\Bigr].$ (4)
Recently, powerful video generative priors based on Diffusion Transformers (DiTs) [peebles2023scalable], including Open-Sora [zheng2024open], Wan2.1 [wan2.1, wan2025], MovieGen [polyak2024movie] have demonstrated unprecedented ability to capture complex spatiotemporal dynamics, paving the way toward world simulators [brooks2024video]. These models achieve exceptional temporal coherence by employing bidirectional attention across all video frames. However, this architectural strength is also a critical bottleneck: the generation of a single frame requires processing the entire sequence, including future frames. This non-causal dependency introduces prohibitive latency and makes them fundamentally incompatible with real-time or streaming applications.
2.2 Diffusion Models for Inverse Problems
Using the diffusion priors defined above to solve inverse problems (Eq. 1) requires combining the pre-trained prior score $s_{\theta}$ with the likelihood $p(\boldsymbol{y}|\boldsymbol{x})$. While this Bayesian framework has achieved great success in static image inverse problems, the field is now moving towards the significantly more challenging video domain, which introduces the critical demand for spatiotemporal consistency.
Diffusion Models for Image Inverse Problems.
In the image domain, sampling-based methods [chung2022diffusion, rout2023solving, song2023solving] work by injecting an approximation of the likelihood gradient, $\nabla_{\mathbf{x}}\log p(\boldsymbol{y}|\boldsymbol{x})$, into the reverse sampling process (Eq. 3). This is often achieved by approximating the intractable likelihood term $\nabla_{\mathbf{x}_{t}}\log p(y|\mathbf{x}_{t})$ using the posterior mean $\hat{\mathbf{x}}_{0}$. Particle-based optimization methods [mardani2023variational, zilberstein2024repulsive] avoid this approximation via variational inference. While effective for producing high-fidelity images, these methods are inherently slow, requiring hundreds or thousands of function evaluations to generate a single result.
Diffusion Models for Video Inverse Problems.
Extending this framework to video inverse problems introduces a significant new challenge: maintaining temporal consistency. The field has primiarily focused on adapting the image-based solvers by adding temporal heuristics. For example, Warped Diffusion [daras2024warped] uses optical flow to warp noise, SVI [kwon2025solving] and Vision-XL [kwon2024visionxlhighdefinitionvideo] treat time as a batch dimension, while STEP [zhang2025step] tuned image diffusion model to capture temporal consistency. These methods are constrained by a weak image-native prior that struggles with complex dynamics and remain prohibitively slow. To address this speed bottleneck, one-step distillation methods [wang2025seedvr2, zhuang2025flashvsr] have been proposed for specific tasks like super-resolution. However, they are highly task-specific and relies on massive, fully-paired $(\boldsymbol{x},\boldsymbol{y})$ datasets (e.g., 10M video pairs), limiting its generality and scalability.
This leaves a clear research gap: existing methods force a trade-off between weak-but-slow image priors, or fast-but-task-specific models that require massive paired data. This motivates our work, InstantViR, which introduces an amortized paradigm to distill powerful, general-purpose video priors (e.g., Wan2.1 [wan2.1]) into a one-step, feed-forward solver that operates without paired data, enabling real-time, high-fidelity reconstruction for the first time.
Figure 2: Overview of the InstantViR framework. (Top) Training: We train a single-step solver $q_{\phi}$ using only degraded measurements $\boldsymbol{y}$. The solver is optimized with two objectives: a data fidelity loss (ensuring the reconstruction matches the measurement $\boldsymbol{y}$) and a prior distillation loss (using a frozen video diffusion prior [wan2.1], $s_{\theta}$, to enforce temporal consistency and realism). (Bottom) Inference: The trained solver $q_{\phi}$ operates as a fast, feed-forward network, processing the video in a causal, block-wise, and autoregressive manner. This enables real-time, single-step streaming reconstruction with optional text guidance.
3 Method
We now describe InstantViR, which transforms a slow, iterative video diffusion prior into a fast, single-step solver for video inverse problems. In Sec. 3.1, we introduce our variational amortized inference formulation and derive an objective that enables training without paired data. In Sec. 3.2, we present our causal video inverse solver architecture, which enables efficient block-wise streaming inference. In Sec. 3.3, we show how to replace the original video VAE with an ultra-efficient LeanVAE while carefully handling the induced latent distribution shift, which is crucial for achieving true real-time performance.
3.1 Amortized Inference Framework
Modern video diffusion models typically operate in a latent space. A video $\boldsymbol{x}$ is first encoded into latent codes $\boldsymbol{z}=\mathcal{E}(\boldsymbol{x})$ by a video VAE, and then a Diffusion Transformer (DiT) backbone [peebles2023scalable] learns a generative prior $p(\boldsymbol{z})$ over these latents. At sampling time, frames are decoded via $\boldsymbol{x}=\mathcal{D}(\boldsymbol{z})$, yielding high-fidelity, temporally coherent videos. Our goal is to leverage such a pre-trained video diffusion model as a prior to solve video inverse problems. Given degraded measurements $\boldsymbol{y}$ related to the clean video $\boldsymbol{x}$ via a known forward operator $\boldsymbol{y}=\mathcal{A}(\boldsymbol{x})$, we wish to sample from the posterior $p(\boldsymbol{x}|\boldsymbol{y})$ defined in Eq. 1.
Traditional sampling-based approaches [daras2024warped, kwon2025solving, kwon2024visionxlhighdefinitionvideo] inject the likelihood $p(\boldsymbol{y}|\boldsymbol{x})$ (defined by $\mathcal{A}$) into the diffusion process and iteratively perturb the generation trajectory to enforce measurement consistency. This yields high quality but is extremely slow, and repeatedly backpropagating through the decoder $\mathcal{D}$ adds further overhead.
Our amortized inference approach, instead, directly learns a single-step feed-forward solver, $q_{\phi}(\boldsymbol{x}|\boldsymbol{y})$, as a variational approximation to the posterior $p(\boldsymbol{x}|\boldsymbol{y})$. We minimize the expected KL divergence
$\mathcal{L}=\mathbb{E}_{\boldsymbol{y}\sim p(\boldsymbol{y})}\Big[D_{\mathrm{KL}}\!\left(q_{\phi}(\boldsymbol{x}|\boldsymbol{y})\,\|\,p(\boldsymbol{x}|\boldsymbol{y})\right)\Big],$ (5)
which, up to an additive constant, can be decomposed as
$\begin{split}\mathbb{E}_{\boldsymbol{y}}\Bigl\{\mathbb{E}_{\boldsymbol{x}\sim q_{\phi}(\boldsymbol{x}|\boldsymbol{y})}\Big[-\log p(\boldsymbol{y}|\boldsymbol{x})\Big]+D_{\mathrm{KL}}\!\left(q_{\phi}(\boldsymbol{x}|\boldsymbol{y})\,\|\,p(\boldsymbol{x})\right)\Bigl\},\end{split}$ (6)
where the first term enforces data fidelity through the known degradation model, and the second term regularizes the solver to remain on the natural video manifold defined by the diffusion prior $p(\boldsymbol{x})$ (e.g., Wan2.1 [wan2.1]).
Because the diffusion prior is defined in the latent space, we work with $\boldsymbol{z}$ instead of $\boldsymbol{x}$, where $\boldsymbol{x}=\mathcal{D}(\boldsymbol{z})$ and $\boldsymbol{z}\sim p(\boldsymbol{z})$. We parameterize the solver as $q_{\phi}(\boldsymbol{z}|\boldsymbol{y})$ and obtain the latent-space objective
$\mathbb{E}_{\boldsymbol{y}}\bigl\{\underbrace{\mathbb{E}_{\boldsymbol{z}\sim q_{\phi}(\boldsymbol{z}|\boldsymbol{y})}\left[-\log p\big(\boldsymbol{y}|\mathcal{D}(\boldsymbol{z})\big)\right]}_{\mathcal{L}_{\text{likelihood}}}+\underbrace{D_{\mathrm{KL}}\!\left(q_{\phi}(\boldsymbol{z}|\boldsymbol{y})\,\|\,p(\boldsymbol{z})\right)}_{\mathcal{L}_{\text{prior}}}\bigl\}.$ (7)
The prior term $\mathcal{L}_{\text{prior}}$ involves a KL divergence to the diffusion prior $p(\boldsymbol{z})$, which is only implicitly defined via its score function $s_{\theta}(\boldsymbol{z}_{t},t)$. Following score-distillation methods [luo2023diff, yin2024one, lee2024diffusion], we approximate this KL by a score-matching loss:
$\mathcal{L}_{\text{prior}}\approx\mathbb{E}_{t,\boldsymbol{\epsilon},\boldsymbol{z}\sim q_{\phi}(\cdot|\boldsymbol{y})}\big[w(t)\,\|s_{\theta}(\boldsymbol{z}_{t},t)-s_{q_{\phi}}(\boldsymbol{z}_{t},t)\|^{2}\big],$ (8)
where $\boldsymbol{z}_{t}=\alpha_{t}\boldsymbol{z}+\sigma_{t}\boldsymbol{\epsilon}$ is the noised latent at time $t$, $w(t)$ is a weighting schedule, and $s_{q_{\phi}}$ denotes the score implied by the solver (implemented via a small auxiliary head $s_{\varphi}$).
Optimizing the combined objective in Eq. 7 trains $q_{\phi}(\boldsymbol{z}|\boldsymbol{y})$ to provide an amortized approximation of the true posterior $p(\boldsymbol{z}|\boldsymbol{y})$, and thus $p(\boldsymbol{x}|\boldsymbol{y})$ via $\mathcal{D}$. This training requires only (i) the degraded measurements $\boldsymbol{y}$ and (ii) the frozen pre-trained diffusion model $s_{\theta}$. No paired $(\boldsymbol{x},\boldsymbol{y})$ ground-truth dataset is needed, making the framework naturally scalable and flexibly conditioned on arbitrary measurement operators.
3.2 Causal Autoregressive Solver
A key distinction between offline and streaming video processing is causality. In offline generation, the model can attend to the entire video sequence, whereas in streaming reconstruction it only has access to past and current frames at time step $n$, not future ones. As a result, a standard DiT with full spatiotemporal attention cannot be directly used as a streaming video inverse solver.
Following this constraint, we design $q_{\phi}$ as a causal autoregressive solver that operates on temporal blocks of $T$ frames (Fig. 2), inspired by [yin2024slow]. Each block of degraded frames $\boldsymbol{y}_{n}$ is encoded to latents, and the model processes these blocks sequentially using a block-causal attention mechanism with two modes:
  • Intra-block bidirectional attention. Within the current block $n$, all tokens attend to each other to model rich local spatiotemporal structure. For a query $\mathbf{Q}_{i}$ from token $i$ in block $n$, with keys $\mathbf{K}_{n}$ and values $\mathbf{V}_{n}$ from all tokens in that block,
    $\mathrm{Att}_{\mathrm{intra}}(\mathbf{Q}_{i},\mathbf{K}_{n},\mathbf{V}_{n})=\mathrm{softmax}\!\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{n}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{n}.$ (9)
  • Inter-block causal attention. Across blocks, tokens in block $n$ are allowed to attend only to previous reconstructed blocks $\hat{\boldsymbol{z}}_{<n}$. For the same query $\mathbf{Q}_{i}$, with keys $\mathbf{K}_{<n}$ and values $\mathbf{V}_{<n}$ from all past blocks,
    $\mathrm{Att}_{\mathrm{inter}}(\mathbf{Q}_{i},\mathbf{K}_{<n},\mathbf{V}_{<n})=\mathrm{softmax}\!\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{<n}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{<n},$ (10)
In practice, we implement the inter-block attention with standard autoregressive KV caching: after reconstructing each block, we store its attention keys and values and reuse them as $\mathbf{K}_{<n}$ and $\mathbf{V}_{<n}$ for subsequent blocks. This avoids redundant computation over past frames while preserving exact causal behavior and keeping the per-frame cost low.
3.3 Latent-Space Adaptation with Compact VAE
The combination of variational diffusion distillation and our causal architecture already yields a strong solver, achieving about 15 FPS at $832\times 480$ resolution with the original video VAE. At this point, however, the DiT backbone is no longer the main bottleneck—the heavy video decoder becomes the limiting factor for real-time performance, consistent with prior observations [wang2025seedvr2, zhuang2025flashvsr] in latent video diffusion models. A natural attempt to accelerate the system would replace the original VAE with a more efficient architecture. However, this breaks the semantics of the latent space: the diffusion prior $p(\boldsymbol{z})$ and score $s_{\theta}$ are trained in the original latent space $\boldsymbol{z}$ induced by $(\mathcal{E},\mathcal{D})$, while a new VAE $(\mathcal{E}^{\prime},\mathcal{D}^{\prime})$ defines a different latent space $\boldsymbol{z}^{\prime}$. Directly distilling a solver in $\boldsymbol{z}^{\prime}$ without accounting for this distribution shift causes severe mismatch with the teacher prior.
To address this, we propose a teacher-space regularized distillation scheme that explicitly bridges the two latent spaces. We train a video inverse solver $q^{\prime}_{\phi}(\boldsymbol{z}^{\prime}|\boldsymbol{y})$ in the $\boldsymbol{z}^{\prime}$-space using the objective
$\begin{split}\mathcal{L}(q^{\prime}_{\phi})=&~\mathbb{E}_{\boldsymbol{y}}\mathbb{E}_{\boldsymbol{z}^{\prime}\sim q^{\prime}_{\phi}(\boldsymbol{z}^{\prime}|\boldsymbol{y})}\big[-\log p\big(\boldsymbol{y}|\mathcal{D}^{\prime}(\boldsymbol{z}^{\prime})\big)\big]\\ &+\mathbb{E}_{t,\boldsymbol{\epsilon},\,\boldsymbol{z}^{\prime}\sim q^{\prime}_{\phi}(\cdot|\boldsymbol{y})}\Big[w(t)\,\big\|s_{\theta}(\boldsymbol{z}_{t},t)-s_{q^{\prime}}(\boldsymbol{z}_{t},t)\big\|^{2}\Big],\end{split}$ (11)
where we map the new latent $\boldsymbol{z}^{\prime}$ back into the teacher latent space via
$\boldsymbol{x}=\mathcal{D}^{\prime}(\boldsymbol{z}^{\prime}),\quad\boldsymbol{z}=\mathcal{E}(\boldsymbol{x}),\quad\boldsymbol{z}_{t}=\alpha_{t}\boldsymbol{z}+\sigma_{t}\boldsymbol{\epsilon}.$
Thus, the likelihood term is evaluated in the new latent space, while the prior term uses the teacher diffusion score $s_{\theta}$ in the original latent space $\boldsymbol{z}$ after passing through $\mathcal{D}^{\prime}$ and $\mathcal{E}$. This constrains the new latent space $\boldsymbol{z}^{\prime}$ so that, when decoded and re-encoded, it remains aligned with the teacher prior, enabling effective one-step distillation under a new VAE.
After exploring several efficient VAEs, we adopt LeanVAE [cheng2025leanvae] as our new video tokenizer. LeanVAE is an ultra-efficient spatiotemporal VAE built on a lightweight NAF backbone with wavelet-based channel compression. Plugging LeanVAE into our teacher-space regularized distillation scheme yields an additional $2\times$ speedup over the original VAE, pushing InstantViR beyond 35 FPS while preserving diffusion-level fidelity and temporal coherence for streaming video inverse problems.
Figure 3: Qualitative comparison for video random inpainting. (Top) On the Open-Sora dataset [lin2024open], our model reconstructs a high-fidelity and temporally consistent video from a 50% masked measurement. The zoomed-in face sequence demonstrates the stability and fine detail of our single-step result. (Bottom) We demonstrate strong zero-shot generalization on the REDS dataset [zhang2018unreasonable], our method generates a sharp, coherent video that is perceptually close to the ground truth.
4 Experiment
We evaluate InstantViR on a range of video inverse problems, compare it against strong diffusion-based baselines, analyze the contribution of its key components (amortized solver, causal architecture, LeanVAE), and demonstrate extended capabilities such as text-guided video reconstruction and editing. Further implementation details and additional results are provided in the supplementary material.
Figure 4: Results for text-guided video reconstruction. InstantViR not only restores high-fidelity details from severely masked measurements (top row) but also uses text prompts to control the reconstruction. (Left) Given a masked input, our method can add specific details like a "black headband" or "wear glasses" as guided by the text. (Right) The framework can generate multi-modal, semantically distinct, and plausible outputs from the same masked input, such as reconstructing the subject with either "close eyes" or "open eyes", demonstrating fine-grained controllability.
4.1 Experimental Setup
Datasets and Base Model.
Our primary training dataset is Open-Sora-v1.1 [lin2024open]. We use 6,000 video clips for training, without using any associated text labels. Following our methodology in Sec. 3.1, we do not require paired $(\mathbf{x},\mathbf{y})$ data: degraded measurements $\mathbf{y}$ are generated on-the-fly from clean videos $\mathbf{x}$ via the known forward operators, and only $\mathbf{y}$ is only used to query the teacher diffusion model. For evaluation, we use a held-out set of 500 Open-Sora videos and the standard REDS30 dataset [nah2019ntire] to test zero-shot generalization. All videos are processed at $832\times 480$ resolution. Our video diffusion prior is the Wan2.1-1.3B model [wan2.1]. Training takes two weeks on 8 NVIDIA A100.
Table 1: Quantitative results of temporal quality and inference speed. The former is evaluated with FVD $\downarrow$ and the latter is with FPS $\uparrow$. Best results are in bold, suboptimal are underlined.
Method FVD $\downarrow$ Avg. FPS $\uparrow$
Inpainting Super-Res. Deblur
DPS 375.81 711.61 783.10 <0.02
DiffIR2VR - 311.61 - 0.12
SVI 219.90 176.60 154.38 0.29
VISION-XL 224.74 172.79 138.79 <0.17
InstantViR (Ours) 136.06 153.13 110.51 13.91
InstantViR (Ours) 132.59 156.43 103.45 35.56
Baselines.
We compare InstantViR against a suite of strong diffusion-based video inverse problem solvers: DPS [chung2022diffusion] as a foundational diffusion posterior sampler; SVI [kwon2025solving] and VISION-XL [kwon2024visionxlhighdefinitionvideo], which adapt image diffusion priors with batch-consistent sampling; and DiffIR2VR [yeh2024diffir2vr], which uses hierarchical latent warping for video restoration. We use official implementations for all baselines to ensure high-quality generation and fair comparisons.
Table 2: Quantitative results of spatial quality. Metrics include PSNR, SSIM, LPIPS. Best results are in bold, suboptimal are underlined.
[…]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
DPS 27.68 0.92 0.32 22.78 0.91 0.46 23.54 0.88 0.46
DiffIR2VR - - - 33.44 0.92 0.33 - - -
SVI 29.42 0.90 0.17 33.85 0.96 0.17 26.93 0.89 0.31
VISION-XL 30.83 0.95 0.25 35.69 0.98 0.24 30.03 0.93 0.28
InstantViR (Ours) 30.54 0.97 0.12 34.91 0.96 0.23 31.85 0.97 0.17
InstantViR (Ours) 31.78 0.96 0.13 27.04 0.95 0.22 31.16 0.97 0.15
Metrics.
We evaluate performance using four complementary criteria: (1) Per-frame reconstruction quality: PSNR and SSIM; (2) Perceptual quality: LPIPS [zhang2018unreasonable]; (3) Temporal consistency: Fréchet Video Distance (FVD) [unterthiner2019fvd], which measures the distributional similarity and temporal coherence across frames; (4) Inference speed: frames per second (FPS) on a single NVIDIA A800 80GB GPU.
4.2 Standard Video Inverse Problems
We first evaluate InstantViR on three canonical video inverse problems: random inpainting, Gaussian deblurring, and $4\times$ super-resolution. Table 2 summarizes per-frame reconstruction and perceptual quality, while Table 1 reports temporal consistency and runtime.
As shown in Table 2, our one-step solver InstantViR (using the original VAE) and its accelerated variant InstantViR (with LeanVAE) achieve state-of-the-art or highly competitive performance across all tasks, surpassing iterative methods such as SVI and VISION-XL in PSNR, SSIM, and LPIPS. This confirms that the amortized student solver successfully inherits the high-fidelity prior of the teacher video diffusion model.
[…]
Random Inpainting.
Fig. 3 shows qualitative results for $50\%$ random inpainting. On the Open-Sora test set, InstantViR recovers fine, high-frequency details (e.g., individual teeth) that are blurred or missing in the baselines. Moreover, the reconstructions remain temporally stable across frames. On the unseen REDS [nah2019ntire] dataset, our method still produces sharp, coherent distant city views, whereas baseline methods often exhibit blur and temporal jitter, demonstrating strong zero-shot generalization.
Gaussian Deblurring.
For Gaussian deblurring (Fig. 5), sampling-based methods such as SVI [kwon2025solving] tend to suffer from noticeable temporal flicker, leading to unstable videos. InstantViR, by contrast, produces sharp and consistent reconstruction across frames, restoring fine details, such as the athlete’s earrings, without sacrificing temporal coherence.
$4\times$ Super-Resolution.
In the $4\times$ super-resolution setting (Fig. 6), InstantViR remains highly competitive with or superior to SVI and VISION-XL. While these baselines are spatially stable, they tend to produce unrealistic jagged edges or overly smoothed details. Our method recovers sharp textures that are both perceptually detailed and temporally aligned with the ground truth.
Time
Mea. (Blur)
Refer to caption Refer to caption Refer to caption
DPS [chung2022diffusion]
Refer to caption Refer to caption Refer to caption
SVI [kwon2025solving]
Refer to caption Refer to caption Refer to caption
Vis-XL [kwon2024visionxlhighdefinitionvideo]
Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption
GT
Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparison for video Gaussian deblurring.
Time
Mea. (LR)
Refer to caption Refer to caption Refer to caption
DiffIR2VR
Refer to caption Refer to caption Refer to caption
SVI [kwon2025solving]
Refer to caption Refer to caption Refer to caption
Vis-XL [kwon2024visionxlhighdefinitionvideo]
Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption
GT
Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative comparison for video super-resolution (4x).
4.3 Text-Guided Video Reconstruction
Because our framework builds on a text-conditional video diffusion prior (Wan2.1), InstantViR and InstantViR can naturally generalize to text-guided video reconstruction and editing by conditioning the amortized inference objective (Eq. 7) on prompts. As illustrated in Fig. 1 and 4, this enables fine-grained semantic control over the reconstructed content.
Given a partially masked input video and a text prompt, the model generates plausible, semantically aligned completions in the missing regions. For the same input sequence, different prompts (e.g., "wearing glasses" vs. "wearing a headband") produce distinct but temporally consistent edits of the subject. In another example, we can control whether a musician "closes eyes" or "keeps eyes open" while playing the piano. These results demonstrate that InstantViR not only reconstructs measurement-consistent content (likelihood term), but can also hallucinate coherent new details guided by text (prior term).
This unique combination of high speed, causal streaming, and prompt-level controllability opens up new possibilities for real-time interactive applications, such as live broadcast enhancement, streaming video restoration, and on-the-fly creative editing.
5 Conclusion
In this paper, we presented InstantViR, a novel amortized inference framework that resolves the core trade-off between reconstruction quality and speed in video inverse problems. By distilling a powerful bidirectional video diffusion prior into a lightweight, one-step causal autoregressive solver, InstantViR achieves both robust temporal consistency and real-time performance. Our framework eliminates the need for paired ground truth, replaces the heavy video VAE with a lightweight alternative via teacher-space regularization, and bypasses the slow, iterative posterior sampler. We demonstrate these gains on various streaming video tasks, attaining state-of-the-art temporal quality while surpassing 35 FPS.
Limitations and Future Directions.
Despite these advances, our accelerated model with latent-space adaptation still slightly lags behind the distilled model using the original VAE in reconstruction quality. This suggests that latent distribution shift remains a limiting factor; jointly fine-tuning the lightweight VAE to better align its latent space with the teacher’s original space may further close this gap.
Beyond natural video restoration, the label-free, amortized nature of our framework makes it well suited for scaling to larger unlabeled video corpora and for deployment in other real-time domains such as medical video enhancement. We view InstantViR as a step toward practical, diffusion-level video reconstruction that is both fast and controllable.